OpenStudy (anonymous):

someone PLEASE HELP me I will reward a medal to who ever helps!! Functions f(x) and g(x) are shown below: f(x) = x^2 g(x) = x^2 − 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)?

OpenStudy (anonymous):

First of all you have to factor g(x) try finding \[x_{1},x_{2}\] and then use the following formula: \[ax^{2}+bx+c=a(x-x_{1})(x-x_{2})\]

OpenStudy (anonymous):

im sorry. how do I find \[x {1},x{2}\]

OpenStudy (anonymous):

\[ax^{2}+bx+c=0\] from the following formula: \[x_{1,2}=\frac{ -b \pm \sqrt{D} }{ 2a }\] where \[D =b^{2}-4ac\]